OpenStudy (anonymous):

find the exact value of the expression. tan 225 degrees

OpenStudy (anonymous):

225 is a third degree angle, so tan is positive there; tan (225) = tan(45) In 3rd quadrant, you subtract 180 from the given angle

OpenStudy (anonymous):

do you know what tan(45) = ?

OpenStudy (anonymous):

its 1

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

so you are telling me that 1 is the answer? if so that is what i got before I posted this question and 1 is not the right answer..

OpenStudy (anonymous):

1 is the answer.

OpenStudy (anonymous):

tan(225) = tan(45) = 1.

OpenStudy (anonymous):

once again I mistype the question without fixing it. the question is 255. I calculated it for 225 and it was 1 you are correct. but 255 degrees is the answer. sorry

OpenStudy (anonymous):

tan(255) = tan(75)

OpenStudy (anonymous):

so take tan of 15 and 60 and add them together right?

OpenStudy (anonymous):

use formula for tan(x+y)

OpenStudy (anonymous):

use x = 45, y = 30

OpenStudy (anonymous):

tan(45 + 30) =

OpenStudy (anonymous):

It is possible to use 15 and 60 right?

OpenStudy (anonymous):

[tan 45 + tan 30]/[1- tan 45 tan 30]

OpenStudy (anonymous):

you dont want tan 15, becuase you dont know tan 15..its not popular angle

OpenStudy (anonymous):

tan 45 = 1; tan 60 = sqrt(3)

OpenStudy (anonymous):

oh i see so just use 45 and 30 to make it easier for ourselves

OpenStudy (anonymous):

sorry...yes, I meant 45 and 30

OpenStudy (anonymous):

tan 45 = 1; tan 30 = sqrt(3)/3

OpenStudy (anonymous):

substitute into formula for tan(x+y)

OpenStudy (anonymous):

so the answer would be -sqrt(3)+2 right?

OpenStudy (anonymous):

nvm after checking multiple times the answer is sqrt(3)+2. thank you for your help!